Assume that sun radiates like a black body of temperature T. Calculate T using Stefan-Boltzmann law. Take Q=5.67 x 10^-8 Wm^-2K^-4 and Lo=4 x 10^26 W.
1
Expert's answer
2019-06-03T10:35:51-0400

Stefan-Boltzmann law states, that the total energy radiated per unit surface area of a black body across all wavelengths per unit time (the black-body radiant emittance) is directly proportional to the fourth power of the black body's thermodynamic temperature T:


"\frac{P}{A}=\sigma {{T}^{4}}"

where "P=4\cdot {{10}^{26}}W"  is the power of Sun, "A=4\pi R_{\odot }^{2}" is the surface area of Sun, "{{R}_{\odot }}=7\cdot {{10}^{8}}m" is the Solar radius. Then we find


"{{T}^{4}}=\frac{P}{4\pi R_{\odot }^{2}\cdot \sigma }"

or


"T=\sqrt[4]{\frac{P}{4\pi R_{\odot }^{2}\cdot \sigma }}"

Substitute known values


"T=\sqrt[4]{\frac{4\cdot {{10}^{26}}W}{4\pi {{\left( 7\cdot {{10}^{8}}m \right)}^{2}}\cdot 5.67\cdot {{10}^{-8}}W\cdot {{m}^{-2}}\cdot {{K}^{-4}}}}\approx 5800K"

So the Sun has a surface temperature of 5800 K.